OpenStudy (anonymous):

A cylindrical tank with diameter 20in. is half filled with water. How much will the water level in the tank rise if you place a metallic ball with radius 4in. in the tank? Round answer to the nearest tenth.

OpenStudy (wolf1728):

Volume of Ball = 4/3 * PI * r^3 Volume = 4/3 * PI * 64 Volume = 268.0825731063 cubic inches Volume of a cylinder = π • r² • height A disk 1 inch thick (imagine a very flat cylinder) and a radius of 10inches has a volume of PI * r² * 1 314.159265 cubic inches This disk of 1 inch thickness and a volume of 314.159265 cubic inches would raise the water level exactly 1 inch. So a ball of volume of 268.0826 cubic inches would raise the water level by (268.0826 / 314.1593) inches = 0.8533333248 inches or rounded to the nearest tenth = 0.9 inches

OpenStudy (anonymous):

Refer to a Mathematica solution attached.

OpenStudy (wolf1728):

It's nice to know I got it right !!